OpenStudy (anonymous):

15. What is the equation that contains the coordinates (8,4) and (2,1)? A. y = 2x - 3 B. y = 1/2x - 1 C. y = 1/2x D. y = -1/2x + 8 I think it's B?

OpenStudy (anonymous):

No it cant be B because when 8 is substitute for x the left sides solves as 4-1 which = 3 and the left side is 4. The way to check is to plug in the numbers, x being the first is the paratheses and y being the second, to the x and y in the equation.

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

first find the slope, then substitute the slope and a point into the point slope equation, then solve for slope intercept form

OpenStudy (anonymous):

y=mx+b y=1/2x+b (2,1) 1=1/2*2+b 1=1+b b=1-1 b=?

OpenStudy (anonymous):

i assume so